# OFFSET: Segmentation-based Focus Shift Revision for Composed Image Retrieval

### v1

@@ 1. Introduction
In recent years, with the development of multimodal learning techniques (Wu et al., 2025; Bi et al., 2024; Liu et al., 2025; Bi et al., 2025; Wu et al., 2025; Huang et al., 2023; Liu et al., 2025; Bi et al., 2025; Tian et al., 2025; Wang et al., 2024; Tang et al., 2022; Huang et al., 2025), there has been a growing interest in Composed Image Retrieval (CIR) (Vo et al., 2019) as a novel image retrieval paradigm (Han et al., 2023; Li et al., 2025). As illustrated in Figure 1(a), the input of CIR is a multimodal query, comprising a reference image and a modification text. The modification text is used to convey the user’s modification request for the reference image, and the model retrieves the target image based on the retrieval intent conveyed by the multimodal query. CIR’s ability to express complex retrieval intent makes it a promising candidate for applications in areas such as information detection (Qiu et al., 2025; Zhang et al., 2024; Xu and Liu, 2025; Wang et al., 2025; Xu et al., 2025; Tang et al., 2022), image editing (Lu et al., 2023; Li et al., 2025; Lu et al., 2024; Gao et al., 2024; Lu et al., 2024), information prediction (Wu et al., 2024; Qiu et al., 2025; Yu et al., 2025; Huang et al., 2024; Qiu et al., 2025; Chen et al., 2024) and object manipulation (Zhou et al., 2024; Huang et al., 2024; Zhou et al., 2025; Huang et al., 2024; Zhou et al., 2025).
@@
Nevertheless, despite the considerable advances made in previous research (Wen et al., 2024; Li et al., 2025; Song et al., 2024), the field of CIR remains in its nascent stages, ignoring the following two phenomena. 1) The visual inhomogeneity. In the open domain, images often exhibit a complex visual composition, comprising both dominant regions that are strongly correlated with the modification requirements (e.g., “dog” and “sled” in the reference image of Figure 1(b)) and irrelevant noise regions (e.g., “tree” in the reference image of Figure 1(b)). However, previous work has failed to take into account the inhomogeneity of dominant and noisy regions. Treating them equally can lead to noise interference during the feature extraction, which in turn may cause a degradation of the query features and inaccurate feature composition results. In the fashion domain, although the image composition is relatively simple, users tend to modify only some regions of the image, which gives rise to the problem of inhomogeneity. 2) The text-priority in multimodal queries. As illustrated in Figure 1(c), based on the reference image alone, there is a tendency for individuals to direct their attention to the visual elements “dog”, “sled” and “person”, while the elements “snow” and “tree” may not be considered. Nevertheless, when the modification text is taken into account, it becomes evident that the “tree” is also an important visual feature, which belongs to the dominant regions. Previous research equally treats the reference image and the modification text and ignores the semantic priority of the modification text during the composition, which may result in a modification focus bias.
@@ 2. Related Work
Composed Image Retrieval. As a variant of image-text retrieval (Chen et al., 2024), this task aims to integrate the semantics of the reference image and the modification text to retrieve a target image, enabling flexible interaction and retrieval. According to the types of backbones used for feature extraction, current CIR approaches can be broadly classified into two categories, i.e., 1) conventional approaches based on traditional feature extraction models and 2) modern approaches utilizing VLP models for feature extraction. The first category typically extracts image and text features via traditional models, such as ResNet and LSTM, and then merges the corresponding multimodal query for retrieval (Vo et al., 2019; Chen et al., 2020; Wen et al., 2021). Conversely, with the development of attention mechanism (Wu et al., 2023; Liu et al., 2025), the second category (Li et al., 2025; Huang et al., 2025; Fu et al., 2025) utilizes vision and language pre-training (VLP) based models like CLIP (Radford et al., 2021) for feature extraction. Notably, CLIP4Cir (Baldrati et al., 2022) first fine-tunes CLIP and utilizes a simple Combiner to achieve excellent performance, demonstrating a significant potential of VLP models. While these models achieve promising performance in CIR, they often overlook significant contextual focus clues during the composition, which shifts the visual focus and introduces irrelevant noise. To address it, our proposed model performs dual focus mapping, thereby enhancing the focus on the matching information between the image and text context.
@@
Semantic segmentation for mask generation. Semantic segmentation (Yuan et al., 2024; Chen et al., 2023; Zhao et al., 2022; Yu et al., 2025; Yuan et al., 2025; Qian et al., 2024; Yuan et al., 2024; Chen et al., 2023; Zhao et al., 2024; Yuan et al., 2024) has become a powerful technology for generating masks in various computer vision tasks. It is typically employed to create accurate object masks, which serve as crucial inputs for downstream tasks such as image editing, scene understanding, and object removal (Tang et al., 2024; Yuan et al., 2024; Tang et al., 2022). Conventionally, TransUNet (Chen et al., 2021) employs a hybrid architecture that combines a visual transformer encoder with a CNN-based decoder. In the NLP domain, DeepLabV3 (Bucher et al., 2019) synthesizes artificial, pixel-wise features for unseen classes based on word2vec label embeddings. Recently, VLP-based models have significantly impacted the field of semantic segmentation with the development of deep learning technology (Tang et al., 2023; Huang et al., 2025; Zeng et al., 2025; Liu et al., 2025; Tang et al., 2024; Zeng et al., 2024). Although the VLP models (e.g., CLIP (Radford et al., 2021)) are not originally designed for segmentation, they have shown excellent zero-shot capabilities in generating object masks. Building on this, methods such as CLIPSeg (Lüddecke and Ecker, 2022) further refine the process by prompt-guided segmentation, enabling more flexible and interactive mask generation based on the CLIP model. This type of approach offers a promising avenue for mask generation in scenarios where solely limited training data is available or where there are unseen classes. Considering the flexible requirements of the CIR task, our model utilizes mask generation based on CLIPSeg to indirectly guide reference images and modification texts to focus on dominant portions.
@@ 3.3.1. Visual Focus Mapping (VFM)
(3) $&\textbf{F}^{l}_{r}=\varPhi_{\mathbb{I}}\left(x_{r}\right),\\ &\check{\textbf{F}}^{l}_{r}=\varPhi_{\mathbb{I}}\left(x_{r}^{s}\right),\\ \end{aligned}\right.$
@@
(4) $&\textbf{F}^{l(s)}_{r}=\operatorname{Cross\>Attention}\left(Q=\check{\textbf{F}}^{l}_{r},\{K,V\}=\textbf{F}^{l}_{r}\right),\\ &\check{\textbf{F}}^{l(s)}_{r}=\operatorname{Cross\>Attention}\left(Q=\textbf{F}^{l}_{r},\{K,V\}=\check{\textbf{F}}^{l}_{r}\right),\\ \end{aligned}\right.$
@@
(5) $&\tilde{\textbf{F}}^{l{\prime}}_{r}=\operatorname{1x1\>Conv}\left(\left[\textbf{F}^{l(s)}_{r},\check{\textbf{F}}^{l(s)}_{r}\right]^{\top}\right)^{\top},\\ &\tilde{\textbf{F}}^{l}_{r}=\operatorname{FC_{\mathbb{I}}}\left(\tilde{\textbf{F}}^{l{\prime}}_{r}\right),\\ \end{aligned}\right.$
@@ 3.3.2. Textual Focus Mapping (TFM)
(7) $&\textbf{F}^{l{\prime}}_{m}=\varPhi_{\mathbb{T}}\left(t_{m}\right),\\ &\textbf{F}^{g}_{m}=\varPhi_{\mathbb{T}}^{g}\left(\textbf{F}^{l{\prime}}_{m}\right),\\ \end{aligned}\right.$
where $\varPhi_{\mathbb{T}}$ and $\varPhi_{\mathbb{T}}^{g}$ denote the penultimate layer and last layer of the CLIP text encoder, respectively. Then, similar to Eqn.((((4)))), we utilize cross attention to semantically interact $\textbf{F}^{g}_{m}$ and $\check{\textbf{F}}^{g}_{r}$ as each other’s Query to obtain the corresponding $\textbf{F}^{g(s)}_{m},\check{\textbf{F}}^{g(s)}_{r}\in\mathbb{R}^{D}$. And then, we use $\operatorname{1x1}$ convolution to further integrate their specifics, i.e. $\tilde{\textbf{F}}^{g{\prime}}_{m}=\operatorname{1x1\>Conv}([\textbf{F}^{g(s)}_{m},\check{\textbf{F}}^{g(s)}_{r}]^{\top})^{\top}$, and concatenate it with the global feature of the modification text and the dominant segmentation (i.e., $\textbf{F}^{g}_{m},\check{\textbf{F}}^{g}_{r}$). The focus-mapped global feature $\tilde{\textbf{F}}^{g}_{m}$ of the modification text is obtained as follows,
@@
where $\tilde{\textbf{F}}^{g}_{m}\in\mathbb{R}^{3\times D}$. For the local feature of the modification text, similar to Eqn.((((5)))), we utilize the fully connected layer $\operatorname{FC_{\mathbb{T}}}$ for dimensional alignment, yielding $\textbf{F}^{l}_{m}\in\mathbb{R}^{S\times D}$, where $S$ is the sequence channel number of modification text. Finally, we align the textual global and local focus mapping results via MG-FP (detailed in the section 3.3.3) and obtain modification focused feature $\tilde{\textbf{F}}_{m}$.
@@ 3.4. Textually Guided Focus Revision
Subsequently, to push the multimodal composed feature $\tilde{\textbf{F}}_{c}$ closer to the target focused feature $\tilde{\textbf{F}}_{t}$, we employ batch-based classification loss (Vo et al., 2019), which is wildly utilized in CIR, formulated as follows,
@@ 4.1.2. Implementation Details
Following previous work (Wen et al., 2024), OFFSET utilizes pre-trained CLIP (ViT-H/14 version) as the backbone. We trained OFFSET via AdamW optimizer with an initial learning rate of $1e$-$4$, while the learning rate for CLIP is set to $1e$-$6$ for better convergence and the batch size is $16$. We empirically set the embedding dimension $D$ = $1024$. Through hyper-parameter tuning, the focus channel number $P$ is set to $4$, and the temperature factor $\tau$ in Eqn.((((12,14)))) is set to $0.1$ for all datasets. The trade-off hyper-parameter $\mu$ is searched via grid search and finally confirmed as $\mu\!=\!0.5$ on FashionIQ and Shoes, while $\mu\!=\!0.8$ for CIRR. All experiments were performed on a single NVIDIA A40 GPU with $48$GB memory and trained $10$ epochs.
@@
To ensure a fair assessment of model performance across different datasets, we adopt standard evaluation protocols following previous works (Wen et al., 2024; Li et al., 2025). The primary metric is Recall@$k$ (abbreviated as R@$k$). For the FashionIQ dataset, we utilize R@$10$, R@$50$, and their category-wise averages for the fair evaluation. For the Shoes dataset, we calculate R@$k$ ($k\!=\!1,10,50$) and their mean value. CIRR assessment included R@$k$ ($k\!=\!1,5,10,50$), Rsubset@$k$ ($k\!=\!1,2,3$), and the average of R@$5$ and Rsubset@$1$.
[…]
@@
MGUR (2024) (ICLR’24) 32.61 61.34 33.23 62.55 41.40 72.51 35.75 65.47
[…]
Prog. Lrn. (2022) (SIGIR’22) 38.18 64.50 48.63 71.54 52.32 76.90 46.38 70.98
FashionSAP (2023) (CVPR’23) 33.71 60.43 41.91 70.93 33.17 61.33 36.26 64.23
FAME-ViL (2023) (CVPR’23) 42.19 67.38 47.64 68.79 50.69 73.07 46.84 69.75
@@
FashionERN (Chen et al., 2024) (AAAI’24) 50.32 71.29 50.15 70.36 56.40 77.21 52.29 72.95
LIMN (Wen et al., 2023) (TPAMI’24) 50.72 74.52 56.08 77.09 60.94 81.85 55.91 77.82
LIMN+ (Wen et al., 2023) (TPAMI’24) 52.11 75.21 57.51 77.92 62.67 82.66 57.43 78.60
@@
[…]
Method R@$1$ R@$10$ R@$50$ Avg
@@
MGUR (2024) (ICLR’24) 18.41 53.63 79.84 50.63
[…]
@@
LIMN (Wen et al., 2023) (TPAMI’24) - 68.20 87.45 -
LIMN+ (Wen et al., 2023) (TPAMI’24) - 68.37 88.07 -
@@
[…]
Method R@$k$ Rsubset@$k$ (R@5+Rsubset@$1$)/$2$
@@
[…]
LF-CLIP (2022) (CVPR’22) 33.59 65.35 77.35 95.21 62.39 81.81 92.02 63.87
CLIP4CIR (2022) (CVPRW’22) 38.53 69.98 81.86 95.93 68.19 85.64 94.17 69.09
TG-CIR (Wen et al., 2023) (ACM MM’23) 45.25 78.29 87.16 97.30 72.84 89.25 95.13 75.57
FashionERN (Chen et al., 2024) (AAAI’24) - 74.77 - 74.93 - - 74.85
LIMN (Wen et al., 2023) (TPAMI’24) 43.64 75.37 85.42 97.04 69.01 86.22 94.19 72.19
LIMN+ (Wen et al., 2023) (TPAMI’24) 43.33 75.41 85.81 97.21 69.28 86.43 94.26 72.35
SADN (Wang et al., 2024) (ACM MM’24) 44.27 78.10 87.71 97.89 72.71 89.33 95.38 75.41

### v2

@@ 1. Introduction
In recent years, with the development of multimodal learning techniques (Wu et al., 2025a; Bi et al., 2024; Liu et al., 2025d; Bi et al., 2025a; Wu et al., 2025b; Huang et al., 2023; Liu et al., 2025a; Bi et al., 2025b; Tian et al., 2025; Wang et al., 2024a; Tang et al., 2022b; Huang et al., 2025a), there has been a growing interest in Composed Image Retrieval (CIR) (Vo et al., 2019a) as a novel image retrieval paradigm (Han et al., 2023b; Li et al., 2025b). As illustrated in Figure 1(a), the input of CIR is a multimodal query, comprising a reference image and a modification text. The modification text is used to convey the user’s modification request for the reference image, and the model retrieves the target image based on the retrieval intent conveyed by the multimodal query. CIR’s ability to express complex retrieval intent makes it a promising candidate for applications in areas such as information detection (Qiu et al., 2025b; Zhang et al., 2024; Xu and Liu, 2025; Wang et al., 2025; Xu et al., 2025; Tang et al., 2022c), image editing (Lu et al., 2023; Li et al., 2025a; Lu et al., 2024b; Gao et al., 2024; Lu et al., 2024a), information prediction (Wu et al., 2024; Qiu et al., 2025c; Yu et al., 2025a; Huang et al., 2024a; Qiu et al., 2025a; Chen et al., 2024c) and object manipulation (Zhou et al., 2024; Huang et al., 2024b; Zhou et al., 2025a; Huang et al., 2024c; Zhou et al., 2025b).
@@
Nevertheless, despite the considerable advances made in previous research (Wen et al., 2024; Li et al., 2025b; Song et al., 2024), the field of CIR remains in its nascent stages, ignoring the following two phenomena. 1) The visual inhomogeneity. In the open domain, images often exhibit a complex visual composition, comprising both dominant regions that are strongly correlated with the modification requirements (e.g., “dog” and “sled” in the reference image of Figure 1(b)) and irrelevant noise regions (e.g., “tree” in the reference image of Figure 1(b)). However, previous work has failed to take into account the inhomogeneity of dominant and noisy regions. Treating them equally can lead to noise interference during the feature extraction, which in turn may cause a degradation of the query features and inaccurate feature composition results. In the fashion domain, although the image composition is relatively simple, users tend to modify only some regions of the image, which gives rise to the problem of inhomogeneity. 2) The text-priority in multimodal queries. As illustrated in Figure 1(c), based on the reference image alone, there is a tendency for individuals to direct their attention to the visual elements “dog”, “sled” and “person”, while the elements “snow” and “tree” may not be considered. Nevertheless, when the modification text is taken into account, it becomes evident that the “tree” is also an important visual feature, which belongs to the dominant regions. Previous research equally treats the reference image and the modification text and ignores the semantic priority of the modification text during the composition, which may result in a modification focus bias.
@@ 2. Related Work
Composed Image Retrieval. As a variant of image-text retrieval (Chen et al., 2024b), this task aims to integrate the semantics of the reference image and the modification text to retrieve a target image, enabling flexible interaction and retrieval. According to the types of backbones used for feature extraction, current CIR approaches can be broadly classified into two categories, i.e., 1) conventional approaches based on traditional feature extraction models and 2) modern approaches utilizing VLP models for feature extraction. The first category typically extracts image and text features via traditional models, such as ResNet and LSTM, and then merges the corresponding multimodal query for retrieval (Vo et al., 2019a; Chen et al., 2020; Wen et al., 2021). Conversely, with the development of attention mechanism (Wu et al., 2023; Liu et al., 2025d), the second category (Li et al., 2025c; Huang et al., 2025c; Fu et al., 2025) utilizes vision and language pre-training (VLP) based models like CLIP (Radford et al., 2021) for feature extraction. Notably, CLIP4Cir (Baldrati et al., 2022a) first fine-tunes CLIP and utilizes a simple Combiner to achieve excellent performance, demonstrating a significant potential of VLP models. While these models achieve promising performance in CIR, they often overlook significant contextual focus clues during the composition, which shifts the visual focus and introduces irrelevant noise. To address it, our proposed model performs dual focus mapping, thereby enhancing the focus on the matching information between the image and text context.
@@
Semantic segmentation for mask generation. Semantic segmentation (Yuan et al., 2024c; Chen et al., 2023b; Zhao et al., 2022a; Yu et al., 2025b; Yuan et al., 2025; Qian et al., 2024; Yuan et al., 2024d; Chen et al., 2023a; Zhao et al., 2024; Yuan et al., 2024a) has become a powerful technology for generating masks in various computer vision tasks. It is typically employed to create accurate object masks, which serve as crucial inputs for downstream tasks such as image editing, scene understanding, and object removal (Tang et al., 2024a; Yuan et al., 2024b; Tang et al., 2022a). Conventionally, TransUNet (Chen et al., 2021) employs a hybrid architecture that combines a visual transformer encoder with a CNN-based decoder. In the NLP domain, DeepLabV3 (Bucher et al., 2019) synthesizes artificial, pixel-wise features for unseen classes based on word2vec label embeddings. Recently, VLP-based models have significantly impacted the field of semantic segmentation with the development of deep learning technology (Tang et al., 2023; Huang et al., 2025b; Zeng et al., 2025; Liu et al., 2025b; Tang et al., 2024b; Zeng et al., 2024). Although the VLP models (e.g., CLIP (Radford et al., 2021)) are not originally designed for segmentation, they have shown excellent zero-shot capabilities in generating object masks. Building on this, methods such as CLIPSeg (Lüddecke and Ecker, 2022) further refine the process by prompt-guided segmentation, enabling more flexible and interactive mask generation based on the CLIP model. This type of approach offers a promising avenue for mask generation in scenarios where solely limited training data is available or where there are unseen classes. Considering the flexible requirements of the CIR task, our model utilizes mask generation based on CLIPSeg to indirectly guide reference images and modification texts to focus on dominant portions.
@@ 3.3.1. Visual Focus Mapping (VFM)
[…]
@@
[…]
@@
[…]
@@ 3.3.2. Textual Focus Mapping (TFM)
[…]
@@
where $\tilde{\textbf{F}}^{g}_{m}\in\mathbb{R}^{3\times D}$. For the local feature of the modification text, similar to Eqn.((5)), we utilize the fully connected layer $\operatorname{FC_{\mathbb{T}}}$ for dimensional alignment, yielding $\textbf{F}^{l}_{m}\in\mathbb{R}^{S\times D}$, where $S$ is the sequence channel number of modification text. Finally, we align the textual global and local focus mapping results via MG-FP (detailed in the section 3.3.3) and obtain modification focused feature $\tilde{\textbf{F}}_{m}$.
@@ 3.4. Textually Guided Focus Revision
Subsequently, to push the multimodal composed feature $\tilde{\textbf{F}}_{c}$ closer to the target focused feature $\tilde{\textbf{F}}_{t}$, we employ batch-based classification loss (Vo et al., 2019b), which is wildly utilized in CIR, formulated as follows,
@@ 4.1.2. Implementation Details
Following previous work (Wen et al., 2024), OFFSET utilizes pre-trained CLIP (ViT-H/14 version) as the backbone. We trained OFFSET via AdamW optimizer with an initial learning rate of 1e1e-44, while the learning rate for CLIP is set to 1e1e-66 for better convergence and the batch size is $16$. We empirically set the embedding dimension $D$ = $1024$. Through hyper-parameter tuning, the focus channel number $P$ is set to $4$, and the temperature factor $\tau$ in Eqn.((12,14)) is set to $0.1$ for all datasets. The trade-off hyper-parameter $\mu$ is searched via grid search and finally confirmed as $\mu\!=\!0.5$ on FashionIQ and Shoes, while $\mu\!=\!0.8$ for CIRR. All experiments were performed on a single NVIDIA A40 GPU with 4848GB memory and trained $10$ epochs.
@@
To ensure a fair assessment of model performance across different datasets, we adopt standard evaluation protocols following previous works (Wen et al., 2024; Li et al., 2025b). The primary metric is Recall@$k$ (abbreviated as R@$k$). For the FashionIQ dataset, we utilize R@$10$, R@$50$, and their category-wise averages for the fair evaluation. For the Shoes dataset, we calculate R@$k$ (k=1,10,50k\!=\!1,10,50) and their mean value. CIRR assessment included R@$k$ (k=1,5,10,50k\!=\!1,5,10,50), Rsubset@$k$ (k=1,2,3k\!=\!1,2,3), and the average of R@$5$ and Rsubset@$1$.
Table 1. Performance comparison on FashionIQ relative to R@kk($\%$). The overall best results are in bold, while the best results over baselines are underlined.
@@
MGUR (2024d) (ICLR’24) 32.61 61.34 33.23 62.55 41.40 72.51 35.75 65.47
VLP Model-Based Methods
Prog. Lrn. (2022b) (SIGIR’22) 38.18 64.50 48.63 71.54 52.32 76.90 46.38 70.98
FashionSAP (2023b) (CVPR’23) 33.71 60.43 41.91 70.93 33.17 61.33 36.26 64.23
FAME-ViL (2023a) (CVPR’23) 42.19 67.38 47.64 68.79 50.69 73.07 46.84 69.75
@@
FashionERN (Chen et al., 2024a) (AAAI’24) 50.32 71.29 50.15 70.36 56.40 77.21 52.29 72.95
LIMN (Wen et al., 2023a) (TPAMI’24) 50.72 74.52 56.08 77.09 60.94 81.85 55.91 77.82
LIMN+ (Wen et al., 2023a) (TPAMI’24) 52.11 75.21 57.51 77.92 62.67 82.66 57.43 78.60
@@
Table 2. Performance comparison on Shoes with respect to R@kk($\%$). The overall best results are in bold, while the best results over baselines are underlined.
Method R@11 R@$10$ R@$50$ Avg
@@
MGUR (2024d) (ICLR’24) 18.41 53.63 79.84 50.63
VLP Model-Based Methods
@@
LIMN (Wen et al., 2023a) (TPAMI’24) - 68.20 87.45 -
LIMN+ (Wen et al., 2023a) (TPAMI’24) - 68.37 88.07 -
@@
Table 3. Performance comparison on CIRR with respect to R@kk($\%$) and Rsubset@kk($\%$). The overall best results are in bold, while the best results over baselines are underlined.
Method R@$k$ Rsubset@$k$ (R@5+Rsubset@11)/22
@@
VLP Model-Based Methods
LF-CLIP (2022b) (CVPR’22) 33.59 65.35 77.35 95.21 62.39 81.81 92.02 63.87
CLIP4CIR (2022a) (CVPRW’22) 38.53 69.98 81.86 95.93 68.19 85.64 94.17 69.09
TG-CIR (Wen et al., 2023b) (ACM MM’23) 45.25 78.29 87.16 97.30 72.84 89.25 95.13 75.57
FashionERN (Chen et al., 2024a) (AAAI’24) - 74.77 - 74.93 - - 74.85
LIMN (Wen et al., 2023a) (TPAMI’24) 43.64 75.37 85.42 97.04 69.01 86.22 94.19 72.19
LIMN+ (Wen et al., 2023a) (TPAMI’24) 43.33 75.41 85.81 97.21 69.28 86.43 94.26 72.35
SADN (Wang et al., 2024b) (ACM MM’24) 44.27 78.10 87.71 97.89 72.71 89.33 95.38 75.41